# Legacy Procurement Practices Shape How U.S. Cities Govern AI: Understanding Government Employees’ Practices, Challenges, and Needs

### v2

@@ 5.3. Sharing (and assuming) ongoing responsibilities of AI governance
Procurement agreements create a unique type of accountability structure: local governments established to protect public values must rely on third-party vendors to develop AI solutions. But the work of AI governance does not end at development, or even deployment. Review and oversight of procured AI technologies requires labor and expertise to conduct evaluations, respond to incidents, train users, and enforce compliance. In this section, we share cities’ experiences working with AI vendors to evaluate system performance to explore cities’ varying perspectives on delegating AI governance work.
In our conversations, we found that vendors often did not provide performance evaluations after deployment. In response, some participants [P7, P10, P15, P18] evaluated models “on their own”. For example, Hana’s department sought to monitor the effectiveness of a procured gunshot detection AI system to ensure it “was continuing to be effective and meet the needs of the city”. However, the vendor neither shared performance evaluations nor offered guidance on how to assess the system. As a result, Hana and her colleagues had to “develop their own metrics” and monitor performance themselves. Some employees embraced the opportunity to conduct independent evaluations. Hugo, for example, valued the city’s ability to shape evaluation criteria both as a safeguard against vendor’s selective reporting and as an opportunity to strengthen local AI expertise. However, not all employees conducted their own evaluations of procured AI systems, and many participants [P2, P3, P4, P5, P8, P13, P17] believed that conducting performance monitoring and audits should be the responsibility of the vendor.
@@ 6. Conclusion
In this paper, we explore how cities’ varying legacy procurement practices shape how public sector AI technologies are selected, assessed, and governed on the ground. We show how cities’ ability to implement AI governance initiatives that prioritize ethical considerations is heavily shaped by the existing procurement infrastructure and power structures that facilitate (or prohibit) cross-organizational oversight over different types of acquisitions. Understanding these present organizational realities of how AI procurement unfolds is a critical first step towards re-imagining interventions that both acknowledge and address governments’ longstanding purchasing practices head-on.
While government employees are still in early stages of adapting their AI procurement practices, our conversations reveal a deeply concerning trend: many AI vendors are not cooperating with employees’ efforts to understand and mitigate AI harms. The implications of vendors’ non-compliance for the rights and freedoms of vulnerable communities cannot be understated as local, state, and federal governments increasingly outsource the critical work of governance to third-party AI systems  (Mulligan and Bamberger, 2019; Liban, 2024; Hickok, 2024; Booth, 2025). Research to support cities in navigating these power differentials, e.g., by conducting their own evaluations or renegotiating existing contracts, is urgently needed. Finally, we encourage the FAccT community to reflect on what intervening within the structures of public procurement affords, and its limitations as a remedy for algorithmic accountability. Our study prompts us to ask: what types of reforms to cities’ legacy purchasing processes might the FAccT community advocate for? What happens when cities’ negotiations with AI vendors fail to be successful? When might we look towards other approaches, such as direct regulation of AI vendors  (Warburton, 2024; Kang, 2024), to protect the public interest?
@@ 7.1. Ethical Considerations
To preserve anonymity of participating employees and cities, we assured interviewees that their participation was voluntary, they could decline to answer interviewer questions, and their responses would be kept anonymous. For sensitive or potentially identifying interview quotes, we exclude participant IDs to preserve anonymity. When appropriate, we use the “x” character to omit exact dollar amounts to preserve confidentiality. To mitigate the risk that participating cities are identified, we limit the amount of detail we provide about cities’ practices.
@@ Appendix A Defining “procurement”
All the definitions of ”public procurement” that we reviewed encountered share fundamental similarities: they all concern the process of bringing in goods and services that are developed externally, to achieve the goals of a public sector entity. They differ, however, in specific components of the process. For example, while the United States’ federal definition emphasizes a “competitive” purchasing process, denoting the exchange of money as part of procurement, some local governments, like New York City, have definitions that are broader, encompassing all functions related to obtaining goods and services whether or not money changes hands (Aquisition.gov, 2024; of Contract Services, 2024).
@@ Appendix B Cities’ AI Use Cases
Appendix C Cities’ emerging AI governance practices (extended)
@@
Appendix D Interview Protocol
@@ Introduction
If it doesn’t come up naturally, we can prompt them to reflect on specific parts of procurement, such as (1) Planning, (2) RFP writing, (3) Evaluating Vendors, (4) Contracting, (5) Designing/Building/Evaluating the AI, and (6) Deployment, and (7) Post-deployment.

### v3

@@ 5.3. Sharing (and assuming) ongoing responsibilities of AI governance
Procurement agreements create a unique type of accountability relationship: governments established to protect public values must rely on vendors to develop AI solutions. But the work of AI governance does not end at development, or even deployment. Review and oversight of procured AI technologies requires labor and expertise to conduct evaluations, respond to incidents, train users, and enforce compliance. In this section, we share cities’ experiences working with AI vendors to evaluate system performance to explore cities’ varying perspectives on delegating AI governance work.
In our conversations, we found that vendors often did not provide performance evaluations after deployment. In response, some participants [P7, P10, P15, P18] evaluated models “on their own”. For example, Hana’s department sought to monitor the effectiveness of a procured gunshot detection AI system to ensure it “was continuing to be effective and meet the needs of the city.” However, the vendor neither shared performance evaluations nor offered guidance on how to assess the system. As a result, Hana and her colleagues had to “develop their own metrics” and monitor performance themselves.
Some employees embraced the opportunity to conduct independent evaluations. Hugo, for example, valued the city’s ability to shape evaluation criteria both as a safeguard against vendor’s selective reporting and as an opportunity to strengthen local AI expertise. However, not all employees conducted their own evaluations of procured AI systems, and many participants [P2, P3, P4, P5, P8, P13, P17] believed that conducting performance monitoring and audits should be the responsibility of the vendor.
@@ 6. Conclusion
In this paper, we shed light on the emerging AI procurement practices of seven U.S. local governments. We illustrate how legacy purchasing norms that apply to all goods and services – school buses, desktop computers, and AI technologies – shape how AI is assessed, selected, and governed on the ground. We summarize key takeaways from our research that contribute to the growing literature on AI procurement:
  1. (1)
    Local governments’ procurement practices vary substantially and meaningfully across jurisdictions. Cities’ legacy procurement practices that are applicable to AI vary along the three dimensions we identified in Sections 4.1, 4.2, and 4.3, with significant implications for AI governance. This matters because past work on AI procurement often paints a simplified, unified view of how local governments’ procurement processes unfold (e.g.,  (Richardson, 2021; Rubenstein, 2021; World Economic Forum, 2020)). While taking a broad view allows one to make widely applicable recommendations, when it comes to AI procurement, the devil is in the details – of cost thresholds, purchasing pathways, and who is (and isn’t) brought in when making purchasing decisions. We call on researchers to undertake further comparative studies that examine how AI procurement norms vary across local and state contexts. By drawing on lessons from their peers, governments can identify actionable pathways to adapt or reform their own procurement practices.
  2. (2)
    AI technologies are most often not acquired via a formal acquisition process (e.g., using a competitive solicitation/RFP). Instead, most AI technologies are acquired via alternative purchasing pathways that often evade traditional oversight mechanisms for government purchasing, such as purchasing cards, free subscriptions, donations, or piggyback contracts. Most efforts to reform governments’ AI procurement practices target the formal solicitation process  (World Economic Forum, 2020; Brauneis and Goodman, 2018; Conti-Cook, [n. d.]; Dotan et al., 2023), and thus are not applicable to most real-world AI acquisitions.
  3. (3)
    While governments are still in early stages of adapting their AI procurement practices, our conversations reveal a deeply concerning trend: many AI vendors are not cooperating with employees’ efforts to understand and mitigate AI harms. The implications of vendors’ non-compliance for the rights and freedoms of vulnerable communities cannot be understated as local, state, and federal governments increasingly outsource the critical work of governance to third-party AI systems  (Mulligan and Bamberger, 2019; Liban, 2024; Hickok, 2024; Booth, 2025). Efforts to support cities in navigating these power differentials, e.g., by conducting their own evaluations, are urgently needed.
Looking forward, we encourage the FAccT community to reflect on what intervening within the structures of public procurement affords, and its limitations as a remedy for algorithmic accountability. Our study prompts us to ask: what types of reforms to cities’ legacy purchasing processes might the FAccT community advocate for? What happens when cities’ negotiations with AI vendors fail to be successful? When might we look towards other approaches, such as direct regulation of AI vendors  (Warburton, 2024; Kang, 2024), or proposals that shift meaningful decision-making power towards members of the public  (Reisman et al., 2018; Moss et al., 2021), to protect the public interest?
@@ 7.1. Ethical Considerations
To preserve anonymity of participating employees and cities, we assured interviewees that their participation was voluntary, they could decline to answer interviewer questions, and their responses would be kept anonymous. For sensitive or potentially identifying interview quotes, we exclude participant IDs to preserve anonymity. When appropriate, we use the “x” character to omit exact dollar amounts to preserve confidentiality. To mitigate the risk that participating cities are identified, we limit the amount of detail we provide about each cities’ practices. Participants’ department and job titles were modified to reduce the risk of re-identification.
Acknowledgements.
We thank the municipal employees who participated and made this research possible. We thank members of the San José Government AI Coalition and MetroLab Network for formative conversations that shaped this work. We thank Samantha Dalal, Anna Kawakami, members of the Heidari Lab, attendees of Carnegie Mellon’s Fairness, Ethics, Accountability, and Transparency reading group, and our anonymous reviewers for offering feedback on our manuscript. This project was supported by the Responsible AI Seed Fund at Carnegie Mellon’s Block Center for Technology and Society. NJ and HH acknowledge support from the NSF (IIS2040929 and IIS2229881) and PwC (through the Digital Transformation and Innovation Center at CMU). Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not reflect the views of the National Science Foundation and other funding agencies.
@@ Appendix A Defining “procurement”
All the definitions of ”public procurement” that we reviewed encountered share fundamental similarities: they all concern the process of bringing in goods and services that are developed externally, to achieve the goals of a public sector entity. They differ, however, in specific components of the process. For example, while the United States’ federal definition emphasizes a “competitive” purchasing process, denoting the exchange of money as part of procurement, some local governments, like New York City, have definitions that are broader, encompassing all functions related to obtaining goods and services whether or not money changes hands (Aquisition.gov, 2024; NYC Mayor’s Office of Contract Services, 2024).
@@ Appendix B Cities’ AI Use Cases
Appendix C Timeline of Events
Our research on AI procurement, like previous empirical research, captures a particular and unique “snapshot in time”  (Autio et al., 2023). Below, we provide an abbreviated timeline of key events that occurred while we conducted our research:
  • December 2023: Interviews with local government employees begin.
  • March 2024: The San José Government AI Coalition is publicly launched with a membership of over 150 agencies, releasing their first public policy templates for local governments. All U.S. local governments are invited to join.
    Many (but not all) of our participating governments do decide to join, and consider how they might adopt the coalition’s resources.
  • March 2024: The Biden Administration issues a second RFI for M-24-18: “Advancing the Responsible Acquisition of Artificial Intelligence in Government”. Informally referred to as the “AI M-Memo”, the memo establishes requirements that federal agencies must follow when procuring AI systems, requiring additional impact assessments and governance requirements for “rights- or safety-impacting AI”. The memo also outlines a clear vision to increase agency capacity for responsible AI innovation.
    Several participating employees take inspiration from the M-Memo when developing their own AI governance strategies. Our research group responds to the RFI (link), sharing learnings we’ve had since beginning the project.
  • June 2024: Interviews with local government employees end.
  • October 2024: The Biden administration adopts the AI M-Memo, issuing a series of short-term deadlines for federal agencies to comply.
  • January 2025: The Trump administration revokes the Biden Executive Order 14110 and the AI M-Memo on day one of holding office. Despite the orders being revoked, scholars reflect on the widespread impact their ideas have had for governments worldwide.
When writing our paper, we aimed to balance characterizing the rapidly-evolving AI-specific practices that governments have adopted, while trying to identify which findings felt most evergreen. We theorize that any of the legacy, decades-old purchasing practices discussed in Section 4 will continue to shape how AI procurement occurs for years to come. However, because interviewed cities were in early stages (their first year) of standing up new AI governance practices like AI-specific risk assessments (Section 4.4), we acknowledge many of these practices and their implications for AI vendors’ compliance are likely to change as cities continue to iterate and share best practices. Our hope is to surface timely opportunities for the FAccT community to contribute to shaping the future of this rapidly evolving landscape.
Appendix D Cities’ emerging AI governance practices (extended)
@@
Appendix E Interview Protocol
@@ E.2. Walk-through
If it doesn’t come up naturally, we can prompt the participant to reflect on specific parts of procurement, such as (1) Planning, (2) RFP writing, (3) Evaluating Vendors, (4) Contracting, (5) Designing/Building/Evaluating the AI, and (6) Deployment, and (7) Post-deployment.